OpenStudy (anonymous):

simplify the radical expression 500see drawing

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

what factors of 500 are also perfect squares?

OpenStudy (anonymous):

22.36

OpenStudy (anonymous):

100 and 5?

jimthompson5910 (jim_thompson5910):

100 is a perfect square, good 5 is not a perfect square

jimthompson5910 (jim_thompson5910):

but \[\Large \sqrt{500} = \sqrt{100*5}\] \[\Large \sqrt{500} = \sqrt{100}*\sqrt{5}\] \[\Large \sqrt{500} = 10*\sqrt{5}\] hopefully you see how I did all that?

OpenStudy (anonymous):

50 and 10?

jimthompson5910 (jim_thompson5910):

the reason we factored 500 into 100*5 is to take advantage of the fact that \[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\]

OpenStudy (anonymous):

oh ok so the answer would be this

jimthompson5910 (jim_thompson5910):

only one factor needs to be a perfect square (not both)

OpenStudy (anonymous):